the length in centimeters of two sides of a rectangle are consecutive odd integers. the perimeter is 120 cm. what is the area of th rectangle
Answer by checkley77(12844)

x(x+2)=120
x^2+2x-120=0
(x+12)(x-10)=0
x-10=0
x=10 for one side.
10=2=12 for the other side.
Proof:
10*12=120
120-120